OpenStudy (anonymous):

36-4(2)(8) what is the real and complex roots? PLEASE HELP i will medal!!!!

zepdrix (zepdrix):

This looks like a simple algebraic expression involving subtraction and multiplication. What roots?...

zepdrix (zepdrix):

Oh oh oh, ok ok I see what this is :) lol

zepdrix (zepdrix):

So this is the value of your `discriminant`.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

You're dealing with something like this:\[\large\rm \sqrt{36-4(2)(8)}\]And they want you to just look at this part of it,\[\large\rm 36-4(2)(8)\]If it's negative, then you have a negative number under the square root, which corresponds to an imaginary number (complex roots)

zepdrix (zepdrix):

So do the algebra real quick, multiply, then subtract. Do you get negative, positive, or zero as an answer?

OpenStudy (anonymous):

i get -28

zepdrix (zepdrix):

Ok good :) Negative number: So when you take the root of negative,\[\large\rm \sqrt{-28}=i \sqrt{28}\]it gives an imaginary number. So we would have `zero real roots`, and `two complex roots` because complex roots always come in pairs of two. And this is a `quadratic`, which has only two roots.

OpenStudy (anonymous):

okay that makes alot more sense! thank you!!!

zepdrix (zepdrix):

np